Suppose that the weight (in pounds) of an airplane is a linear function of the total amount of fuel (in gallons) in its tank. When graphed, the function gives a line with a slope of 6.
With 53 gallons of fuel in its tank, the airplane has a weight of 2318 pounds. What is the weight of the plane with 28 gallons of fuel in its tank?

Answer by ankor@dixie-net.com(22740)

Suppose that the weight (in pounds) of an airplane is a linear function of the
total amount of fuel (in gallons) in its tank.
When graphed, the function gives a line with a slope of 6.
With 53 gallons of fuel in its tank, the airplane has a weight of 2318 pounds.
:
Use the point/slope formula y - y1 = m(x - x1): where
x1 = 53
y1 = 2318
m = 6
:
y - 2318 = 6(x - 53)
y - 2318 = 6x - 318
y = 6x - 318 + 2318
y = 6x + 2000
:
What is the weight of the plane with 28 gallons of fuel in its tank?
x = 28
y = 6(28) + 2000
y = 168 + 2000
y = 2168 lb with 28 gal of fuel